what is the boiling point of solution prepared by adding 85.0g of solution acetate (nach3co 0) to 375 ml of water
1
Expert's answer
2021-10-10T10:25:09-0400

Molarity of NaCH3COO


=(85÷82) / 0.375 Kg

=2.76 mol/ Kg


Bp constant of water = 0.52°C/m


"\Delta"t= iKbm

=2 ×0.52 ×2.76

=2.8704

B.P = 100+ 2.8704


=102.8704°C